# MonoForce: Learnable Image-conditioned Physics Engine

### v3

@@ IV-E Navigation in Unstructured Environments
where $\alpha$ and $\beta$ are the hyperparameters for the trajectory and waypoint costs, respectively. With the described navigation approach, the robot is able to autonomously navigate in the forest environment with uneven terrain, avoiding obstacles and following the set of waypoints, Figure 17. The Figure 16 shows the constructed point cloud map of the forest environment during the navigation experiment. It also contains an aerial photo of the forest area with the robot’s trajectory and the waypoints. The robot was able to autonomously navigate in the forest environment and traverse the 260-meter-long path.
@@ V Conclusion
In this work, we have presented MonoForce, an explainable, physics-informed, and end-to-end differentiable model that predicts the robot’s trajectory from monocular camera images. As a grey-box model (physics-based and data-driven), it benefits from the end-to-end trainability in different domains while it still retains the determinism of its physics engine and explainability of its estimates of terrain properties. Thanks to the method’s end-to-end differentiability, it is possible to incorporate external sensor measurements in the training pipeline. Additionally, due to the inclusion of the physics information, our model is able not only to generalize better (predict more accurate trajectories than data-driven baselines) but also provide interpretable intermediate outputs, for example friction and robot-terrain interaction forces. The training process is self-supervised; it only requires monocular camera images, lidar scans, and SLAM-reconstructed trajectories. The model learns to recognize non-rigid obstacles similar to those that have been driven over in the reference trajectories while keeping the understanding and prediction capabilities for obstacles it has not encountered yet. It treats them as rigid obstacles until future data prove otherwise.
@@ Acknowledgments
This work was co-funded by the Grant Agency of the CTU in Prague under Project SGS24/096/OHK3/2T/13, the European Union under the project Robotics and advanced industrial production (reg. no. CZ.02.01.01/00/22_008/0004590), the Czech Science Foundation under Project 24-12360S.

### v4

@@ IV-E Autonomous Navigation in Unstructured Environments
where $\alpha$ and $\beta$ are the hyperparameters for the trajectory and waypoint costs, respectively. With the described navigation approach, the robot is able to autonomously navigate in the forest environment with uneven terrain, avoiding obstacles and following the set of waypoints, Fig. 15. The Fig. 14 shows the constructed point cloud map of the forest environment during the navigation experiment. It also contains an aerial photo of the forest area with the robot’s trajectory and the waypoints. The robot was able to autonomously navigate in the forest environment and traverse the 260-meter-long path.
@@ V Conclusion
The model also provides interpretable intermediate outputs, such as friction or contacts, which can serve as an additional source of supervision. If such quantities are not directly measurable, it may still be possible to construct PINN-like losses to provide some further regularization.
The training process is self-supervised; it only requires monocular camera images, lidar scans, and SLAM-reconstructed trajectories. In order to speed up the training process, we also leverage the foundation model for semantic segmentation that delivers rigid classes such as stones or tree trunks.
@@ Acknowledgments
This work was co-funded by the Grant Agency of the CTU in Prague under Project SGS24/096/OHK3/2T/13, the European Union under the project Robotics and advanced industrial production (Roboprox reg. no. CZ.02.01.01/00/22_008/0004590), the Czech Science Foundation under Project 24-12360S.